OpenStudy (anonymous):

HELP PLEASE!!!!!! Find the domain of y = x + 5 if the range is {-4, -1, 0, 3}. Show all of your work.

OpenStudy (whpalmer4):

Work backwards. The domain is the values of \(x\), the range is the values of \(y\). For example, if the range contained the value 6 (it doesn't, this is just an example), then you would do \[y = x+5\]\[6=x+5\]\[6-5 = x+5-5\]\[1 = x\]and include \(1\) in the set of values that make up the domain. Do this process for each member of the range to find all the members of the domain.